OpenStudy (anonymous):

3=(3x^2y^0z^-4)^-3 I cant have any Zero, negative or fraction exponents.

OpenStudy (anonymous):

you do not mean 3= do you?

OpenStudy (anonymous):

Yeah ignore that

OpenStudy (anonymous):

Typo

OpenStudy (anonymous):

\[(3x^2y^0z^{-4})^{-3}\]

OpenStudy (anonymous):

Thats exacty how i have it

OpenStudy (anonymous):

multiply all the exponents by -3, ignore the \[y^0\] because it is 1

OpenStudy (anonymous):

\[3^{-3}x^{-6}z^{12}\] ones with negative exponents go down, rest stays up. \[\frac{z^{12}}{3^3x^6}\]

OpenStudy (anonymous):

trick is to not forget the 3. it gets raised to the -3 as well.

OpenStudy (anonymous):

Basic multiplication rules apply? two negatives = posative, + and - = negative?

OpenStudy (anonymous):

*when multiplying exponents

OpenStudy (anonymous):

of course. all normal rules of arithmetic work even when working with exponents.

OpenStudy (anonymous):

One problem I think you did it wrong

OpenStudy (anonymous):

Wait, im confused. How did oyu get 3^-3

OpenStudy (anonymous):

aha i knew it. you have to raise EVERYTHING to the power of -3 because everything was in the parentheses

OpenStudy (anonymous):

Ahah!

OpenStudy (anonymous):

example \[(2x^3y^{-2})^{-2}\] \[=2^{-2}x^{-6}y^4\] dont forget the 2!

OpenStudy (anonymous):

What about X^3 times X^-5

OpenStudy (anonymous):

is it X^-15?

OpenStudy (anonymous):

or X^2

OpenStudy (anonymous):

*X^-2

OpenStudy (anonymous):

\[x^3x^{-5}=x^{3-5}=x^{-2}=\frac{1}{x^2}\]

OpenStudy (anonymous):

dont forget \[x^{-5} = \frac{1}{x^5}\] so \[x^3x^{-5}=\frac{x^3}{x^5}=\frac{1}{x^2}\]\]

OpenStudy (anonymous):

in other words it makes sense to add the exponents when you multiply whether they are positive or negative.

OpenStudy (anonymous):

Only multiply when its an exponent to another power right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[b^n b^m = b^{n+m}\] \[(b^n)^m=b^{mn}\]

OpenStudy (anonymous):

(-4xy^-5)/(24x^-3) How do i take the recripical of this one?

OpenStudy (anonymous):

\[\frac{-4xy^{-5}}{24x^{-3}}\]

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

if the exponent is negative move it from down to up or from up to down. do not mess with the coefficients. they have no exponents attached.

OpenStudy (anonymous):

\[\frac{-4xx^3}{24y^5}\]

OpenStudy (anonymous):

now all the exponents are positive. of course you can simplify this: \[-\frac{x^4}{6y^5}\]

OpenStudy (anonymous):

thanks dude